OpenStudy (anonymous):

write the partial fraction decomposition of the rational expression please. x/(x-5)(x-6) Thank you!

OpenStudy (tkhunny):

\(\dfrac{A}{x-5} + \dfrac{B}{x-6}\) Go!

OpenStudy (anonymous):

so would the answer? be A/x+6+B/x+5

OpenStudy (tkhunny):

No, two things wrong with that. 1) You have written \(\dfrac{A}{x} + 6 + \dfrac{B}{x} + 5\). That is absolutely not what I wrote. Please remember your order of operations and use parentheses where necessary. 2) You must find the values for A and B.

OpenStudy (anonymous):

sorry I meant to write A/(x+6) + B/(x+5). I am confused on how to go on to solve for A or B. I think I confused myself with all the videos I was watching.

OpenStudy (tkhunny):

No worries. Writing in-line is a bit of an adjustment. It takes more parentheses. You have two fractions. Add them - common denominator and that sort of thing.

OpenStudy (anonymous):

So x=A(x+6)+B(x+5)?

OpenStudy (tkhunny):

How did we get "+" instead if "-"? Other than that, yes.

OpenStudy (anonymous):

your right. I have it on my paper as - but for some reason I copied it as + on here. So B turns out to be -5 and x turns out to be -6?